Formulate a system of equations that represents a situation in a school. Describe the method that you would use to solve the system. Then solve the system and explain what the solution means. Thanks in advanced!

Answer by fractalier(6550)

There are 200 students in a school. There are 20 more boys than girls. Find the number of girls and boys in the school
Let b and g be the number of boys and girls.
Thus we have
b + g = 200 and
b = g + 20
Now substitute the second equation into the first and get
g+20 + g = 200
2g + 20 = 200
2g = 180
g = 90 girls
b = g + 20 = 110 boys